### v1

# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

Abstract
Training machine learning models on decentralized private data via federated learning (FL) poses two key challenges: communication efficiency and privacy protection. In this work, we address these challenges within the trusted aggregator model by introducing a novel approach called the Communication-Efficient and Privacy-Adaptable Mechanism (CEPAM), achieving both objectives simultaneously. In particular, CEPAM leverages the rejection-sampled universal quantizer (RSUQ), a construction of randomized vector quantizer whose resulting distortion is equivalent to a prescribed noise, such as Gaussian or Laplace noise, enabling joint differential privacy and compression. Moreover, we analyze the trade-offs among user privacy, global utility, and transmission rate of CEPAM by defining appropriate metrics for FL with differential privacy and compression. Our CEPAM provides the additional benefit of privacy adaptability, allowing clients and the server to customize privacy protection based on required accuracy and protection. We assess CEPAM’s utility performance using MNIST dataset, demonstrating that CEPAM surpasses baseline models in terms of learning accuracy.
I Introduction
Federated learning (FL) [1] enables the training of machine learning models using vast amounts of data held privately by edge devices by exploiting their computational capabilities [2], thanks to its decentralized design [3]. However, FL presents challenges [4], such as ensuring efficient communication during training phases and preserving the privacy of the sensitive data at edge devices. These challenges are commonly addressed separately through methods that introduce some distortion to updated models, like compression methods [5, 6, 7, 8, 9] and differential privacy (DP) mechanisms [10, 11].
To address both challenges simultaneously, a common approach involves initially applying a DP mechanism (such as Laplace [10] or Gaussian [11]) followed by quantization [12, 13]. However, this approach has two drawbacks. Firstly, the overall error (comprising privacy-preserving and quantization errors) will no longer be exact, leading to a degradation in the accuracy of model updates and the final trained model. Secondly, the quantization noise does not contribute to enhancing privacy against the honest but inquisitive clients within the FL framework. In essence, the DP-then-quantize approach is suboptimal.
An alternative approach relies on randomized quantization, such as subtractive dithering [14, 15, 16, 17, 18]. Previous works [19, 20, 21, 22, 23] have employed this method to develop a joint privacy and quantization mechanism for FL. However, these approaches are limited to scalar quantization, which is less effective than vector quantization in terms of compression ratio. Moreover, Lang et al. [24] investigated the use of subtractive dithering to generate privacy-preserving noises such as multivariate $t$-distributed and Laplace noises. However, their proposed mechanism did not encompass the Gaussian mechanism.
Channel simulation [25, 26, 27, 28, 29, 30] is technique that can achieve the same goal. Recognizing that the Gaussian distribution can be represented as a mixture of uniform distributions [31, 32], Agustsson et al. [33] combined subtractive dithering with layered construction [34] to simulate 1D Gaussian channel. This approach further developed in [35] to simulate any 1D unimodal additive noise distribution using randomized scaling and offsets. Recently, a new construction in [36] named RSUQ, which is a randomized vector quantization, was introduced to exactly simulate any multivariate nonuniform continuous additive noise distribution with finite communication by extending 1D methods to multivariate distributions. However, [36] did not consider FL setup. A closely related DP mechanism compressor, based on Poisson functional representation [27], was observed in [37].
In this work, we propose a joint mechanism, CEPAM, aimed at enhancing privacy while compressing model updates in FL. Our proposed CEPAM leverages RSUQ [36], building upon its capability to convert quantization distortion into an additive noise term with adjustable variance, independent of the quantized data. Within CEPAM, both clients and the server can tailor the privacy protection mechanism, offering the choice between the exact Gaussian mechanism (CEPAM-Gaussian) or the exact Laplace mechanism (CEPAM-Laplace), based on the required accuracy level and privacy protection. Additionally, the inherent nature of RSUQ as a vector quantizer gives CEPAM the advantage in reducing the compression ratio compared to its scalar counterparts [22, 23, 33, 34, 35].
We theoretically analyze the performance of CEPAM by studying trade-offs between privacy, rate and accuracy. Specifically, we examine the FL convergence of stochastic gradient descent (SGD)-based federated averaging (FedAvg) with CEPAM in a statistically heterogeneous setup [1, 4], where the dataset available at each user follows different distributions. Additionally, we analyze the privacy of CEPAM-Gaussian and CEPAM-Laplace using techniques from privacy amplification [38, 39]. We validate our theoretical results by comparing CEPAM against other baselines using the MNIST dataset with MLP and CNN architectures. Experimental results demonstrate that CEPAM-Gaussian achieves a 0.5-2.0% improvement in accuracy compared to other baselines.
I-A Related Works
While the original FedAvg [1] works even when the data are non-iid, Stich [40] proposed Local SGD by assuming that the data is iid and provided a theoretical analysis. Later, Li et al. [41] extended Stich’s Local SGD to the heterogeneous data setup and provided a theoretical analysis. Moreover, federated learning with quantization was considered in [8, 9].
Notations
Write $H(X)$ for the entropy and in bits. Logarithms are to the base $2$. For $\mathcal{A},\mathcal{B}\subseteq\mathbb{R}^{n}$, $\beta\in\mathbb{R}$, $\mathbf{G}\in\mathbb{R}^{n\times n}$, $\mathbf{x}\in\mathbb{R}^{n}$ write $\beta\mathcal{A}:=\{\beta\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathbf{G}\mathcal{A}:=\{\mathbf{G}\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathbf{x}:=\{\mathbf{z}+\mathbf{x}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathcal{B}=\{\mathbf{y}+\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$ for the Minkowski sum, $\mathcal{A}-\mathcal{B}=\{\mathbf{y}-\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$, and $\mu(\mathcal{A})$ for the Lebesgue measure of $\mathcal{A}$. Let $B^{n}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}\|\leq 1\}$ be the unit $n$-ball.
II System Model and Preliminaries
In this section, we describe the setup of FL framework. We begin by reviewing the conventional FL framework with bit-constrained model updates in Section II-A, followed by an exploration of LRSUQ in Section II-B, and a discussion of differential privacy and privacy amplification in Section II-C. Next, we formulate the problem and identify the desired properties used in the FL framework in Section II-D.
II-A Federated Learning (FL)
In this work, we consider the FL framework (or federated optimization) proposed in [1]. More explicitly, $K$ clients (or devices), each attached with a local dataset $\mathcal{D}^{(k)}$ where $k\in\{1,2,...,K\}=:\mathcal{K}$, cooperate together to train a shared global model $\mathbf{W}$ with $m$ parameters through a central server. The goal is to minimize the objective function $F:\mathbb{R}^{m}\to\mathbb{R}$:
$\min_{\mathbf{W}\in\mathbb{R}^{m}}\Big{\{}F(\mathbf{W}):=\sum_{k\in\mathcal{K}}p_{k}F_{k}(\mathbf{W})\Big{\}},$ (1)
where $p_{k}$ is the weight of client $k$ such that $p_{k}\geq 0$ and $\sum_{k\in\mathcal{K}}p_{k}=1$. Suppose that the $k$-th local dataset contains $n_{k}$ training data: $\mathcal{D}^{(k)}=\{\xi_{k,1},\dots,\xi_{k,n_{k}}\}$. The local objective function $F_{k}:\mathbb{R}^{m}\to\mathbb{R}$ is defined by
$F_{k}(\mathbf{W}):=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\ell(\mathbf{W};\xi_{k,j}),$ (2)
where $\ell(\cdot;\cdot)$ is an application-specified loss function.
Let $T$ denote the total number of iterations in FL and let $\mathcal{T}_{T}:=\{0,\tau,2\tau,\ldots,T\}$ denote the set of integer multiples of some positive integer $\tau\in\mathbb{Z}^{+}$ where $T\equiv 0\pmod{\tau}$, called the set of synchronization indices. We describe one FL round of the conventional FedAvg [1] for solving optimization problem (1) as follows. Let $\mathbf{W}_{t}$ denote global parameter vector available at the server at the time instance $t\in\mathcal{T}_{T}$. At the beginning of each FL round, the server broadcasts $\mathbf{W}_{t}$ to all the clients. Then, each client $k$ sets $\mathbf{W}_{t}^{k}=\mathbf{W}_{t}$ and computes the $\tau\;(\geq 1)$ local parameter vectors by SGD: [^1] When $T$ is fixed, the larger $\tau$ is, the fewer the communication rounds.
$\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t}^{k}}(\mathbf{W}_{t}^{k}),\;\;t^{\prime}=1,\ldots,\tau,$ (3)
where $\eta_{t+t^{\prime}}$ is the learning rate, $\nabla F_{k}^{j}(\mathbf{W}):=\nabla F_{k}(\mathbf{W};\xi_{k,j})$ is the gradient computed at a single sample of index $j$, and $j_{t}^{k}$ is the sample index chosen uniformly from the local data $\mathcal{D}^{(k)}$ of client $k$ at time $t$222 In this work, our focus is on analyzing the computation of a single stochastic gradient at each client during every time instance. The FL convergence rates can potentially be enhanced by incorporating mini-batching techniques [40], leaving the detailed analysis for future work. Finally, by assuming that all clients participate in each FL round for simplicity, the server aggregates the $K$ local model updates $\{\mathbf{X}_{t+\tau}^{k}=\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}\}_{k\in\mathcal{K}}$ and computes the new global parameter vector:
$\mathbf{W}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\mathbf{X}_{t+\tau}^{k}=\sum_{k\in\mathcal{K}}p_{k}\mathbf{W}_{t+\tau}^{k}.$ (4)
By an abuse of notation, suppose the data in client $k$ is iid sampled from the distribution $\mathcal{D}^{(k)}$. Consequently, the overall distribution becomes a mixture of all local distributions: $\mathcal{D}=\sum_{k\in\mathcal{K}}p_{k}\mathcal{D}^{(k)}$. Previous works typically assume that the data is iid generated by or partitioned among the $K$ clients, i.e., for all $k\in\mathcal{K}$, $\mathcal{D}^{(k)}=\mathcal{D}$. However, we consider a scenario where the data is non-iid (or heterogeneous), implying that $F_{k}$ could potentially be an arbitrarily poor approximation to $F$.
In [9], Shlezinger et al. introduced a framework for Quantized Federated Learning. Given the limited bandwidth of uplink channel in FL framework, each $k$-th client is required to communicate a quantized version [42] of the model update $\mathbf{X}_{t+\tau}^{k}$ using a finite number of bits to the server. Specifically, the $k$-th model update $\mathbf{X}_{t+\tau}^{k}$ is encoded into a binary codeword $u_{t+\tau}^{k}$ of length $R_{k}$ bits through an encoding function $\mathrm{Enc}_{t+\tau}^{k}:\mathbb{R}^{m}\to\{0,1,\ldots,2^{R_{k}}-1\}:=\mathcal{U}_{k}$. Since the outputs $\mathrm{Enc}_{t+\tau}^{k}(\mathbf{X}_{t+\tau}^{k})=u_{t+\tau}^{k}\in\mathcal{U}_{k}$ of an encoder may have unequal probabilities, entropy coding [43, 44, 45] can be utilized to further reduce redundancy when transmitting through the binary lossless channel. In the FL literature, the uplink channel is typically modeled as a bit-constrained link [9], and the transmission is assumed to be lossless. Upon receiving the set of codewords $\{u_{t+\tau}^{k}\}_{k\in\mathcal{K}}$ from clients, the server uses the joint decoding function $\mathrm{Dec}_{t+\tau}:\mathcal{U}_{1}\times\ldots\times\mathcal{U}_{K}\to\mathbf{R}^{m}$ to reconstruct $\hat{\mathbf{X}}_{t+\tau}\in\mathbb{R}^{m}$, an estimate of the weighted average $\sum_{k\in\mathcal{K}}p_{k}\mathbf{X}_{t+\tau}^{k}$.
II-B Rejection-Sampled Universal Quantizer (RSUQ)
We will now review RSUQ [36], which is constructed based on the subtractive dithered quantizer (SDQ) [14, 46, 47].
Given a non-singular generator matrix $\mathbf{G}\in\mathbb{R}^{n\times n}$, a lattice is the set $\mathbf{G}\mathbb{Z}^{n}=\{\mathbf{G}\mathbf{j}:\,\mathbf{j}\in\mathbb{Z}^{n}\}$. A bounded set $\mathcal{P}\subseteq\mathbb{R}^{n}$ is called a basic cell of the lattice $\mathbf{G}\mathbb{Z}^{n}$ if $(\mathcal{P}+\mathbf{G}\mathbf{j})_{\mathbf{j}\in\mathbb{Z}^{n}}$ forms a partition of $\mathbb{R}^{n}$ [48, 49]. Specifically, the Voronoi cell $\mathcal{V}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\arg\min_{\mathbf{j}\in\mathbb{Z}^{n}}\|\mathbf{x}-\mathbf{G}\mathbf{j}\|=\mathbf{0}\}$ is a basic cell. Given a basic cell $\mathcal{P}$, we can define a lattice quantizer $Q_{\mathcal{P}}:\mathbb{R}^{n}\to\mathbf{G}\mathbb{Z}^{n}$ such that $Q_{\mathcal{P}}(\mathbf{x})=\mathbf{y}$ where $\mathbf{y}\in\mathbf{G}\mathbb{Z}^{n}$ is the unique lattice point that satisfies $\mathbf{x}\in-\mathcal{P}+\mathbf{y}$. The resulting quantization error $\mathbf{z}:=Q_{\mathcal{P}}(\mathbf{x})-\mathbf{x}$ depends deterministically on the input $\mathbf{x}$ and is approximately uniformly distributed over the basic cell of the lattice quantizer under some regularity assumptions. Therefore, it is often combined with probabilistic methods such as random dithering to construct SDQ [46, 47].
Definition 1
Given a basic cell $\mathcal{P}$ and a random dither $\mathbf{V}\sim\mathrm{Unif}(\mathcal{P})$, a subtractive dithered quantizer (SDQ) $Q_{\mathcal{P}}^{SDQ}:\mathbb{R}^{n}\times\mathcal{P}\to\mathbb{R}^{n}$ for an input $\mathbf{x}\in\mathbb{R}^{n}$ is given by $Q_{\mathcal{P}}^{SDQ}(\mathbf{x},\mathbf{v})=Q_{\mathcal{P}}(\mathbf{x}-\mathbf{v})+\mathbf{v}$, where $Q_{\mathcal{P}}$ is the lattice quantizer.
A well-known property of SDQ is that the resulting quantization error is uniformly distributed over the basic cell of the quantizer and is statistically independent of the input signal [15, 47, 50, 51, 52].
The quantization error incurred from the aforementioned quantization schemes approximately follows a uniform distribution over a basic cell of a lattice. It may be desirable to have the quantization error follow a uniform distribution over an arbitrary set, rather than being distributed uniformly over a basic cell. RSUQ is a randomized quantizer where the quantization error is uniformly distributed over a set $\mathcal{A}$, a subset of a basic cell. This quantization scheme is based on applying rejection sampling on top of SDQ. Intuitively, we keep generating new dither signals until the quantization error falls in $\mathcal{A}$.
Definition 2
[36, Definition 4] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a subset $\mathcal{A}\subseteq\mathcal{P}$, and a sequence $S=(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}}$, $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ are i.i.d. dithers, the rejection-sampled universal quantizer (RSUQ) $Q_{\mathcal{A},\mathcal{P}}:\mathbb{R}^{n}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $\mathcal{A}$ against $\mathcal{P}$ is given by
$Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},(\mathbf{v}_{i})_{i}):=Q_{\mathcal{P}}(\mathbf{x}-\mathbf{v}_{h})+\mathbf{v}_{h},$ (5)
where
$h:=\min\big{\{}i:\,Q_{\mathcal{P}}(\mathbf{x}-\mathbf{v}_{i})+\mathbf{v}_{i}-\mathbf{x}\in\mathcal{A}\big{\}},$ (6)
and $Q_{\mathcal{P}}$ is the lattice quantizer for basic cell $\mathcal{P}$.
Note that SDQ is a special case of RSUQ where $\mathcal{A}=\mathcal{P}$. It is easy to check that the quantization error is uniform over $\mathcal{A}$ by using the standard rejection sampling argument and the “crypto” lemma [49, Lemma 4.1.1 and Theorem 4.1.1].
Proposition 3
[36, Proposition 5] For any $\mathbf{x}\in\mathbb{R}^{n}$, the quantization error $\mathbf{Z}:=Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},S)-\mathbf{x}$ of RSUQ $Q_{\mathcal{A},\mathcal{P}}$, where $S\sim P_{S}$, follows the uniform distribution over the set $\mathcal{A}$, i.e., $\mathbf{Z}\sim\mathrm{Unif}(\mathcal{A})$.
By using layered construction as in [35, 34], RSUQ can be generalized to simulate an additive noise channel with noise following a continuous distribution.
Definition 4
[36, Definition 12] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a probability density function $f:\mathbb{R}^{n}\to[0,\infty)$ where $L_{u}^{+}(f)$ is always bounded for $u>0$, and $\beta:(0,\infty)\to[0,\infty)$ satisfying $L_{u}^{+}(f)\subseteq\beta(u)\mathcal{P}$ for $u>0$, and a random pair $S=(U,(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}})$ where the latent variable $U\sim f_{U}$ with $f_{U}(u):=\mu(L_{u}^{+}(f))$, and $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ is a sequence of i.i.d. dither signals, the layered rejection-sampled universal quantizer (LRSUQ) $Q_{f,\mathcal{P}}:\mathbb{R}^{n}\times\mathbb{R}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $f$ against $\mathcal{P}$ is given by
$Q_{f,\mathcal{P}}(\mathbf{x},u,(\mathbf{v}_{i})_{i}):=\beta(u)\cdot\big{(}Q_{\mathcal{P}}(\mathbf{x}/\beta(u)-\mathbf{v}_{h})+\mathbf{v}_{h}\big{)},$ (7)
where
$h:=\min\big{\{}i:\,\beta(u)\cdot\big{(}Q_{\mathcal{P}}(\mathbf{x}/\beta(u)-\mathbf{v}_{i})+\mathbf{v}_{i}\big{)}-\mathbf{x}\in L_{u}^{+}(f)\big{\}},$ (8)
and $Q_{\mathcal{P}}$ is the lattice quantizer for basic cell $\mathcal{P}$.
It can be shown that LRSUQ indeed gives the desired error distribution by Proposition 3.
Proposition 5
[36, Proposition 13] Consider LRSUQ $Q_{f,\mathcal{P}}$. For any random input $\mathbf{X}$, the quantization error defined by $\mathbf{Z}:=Q_{f,\mathcal{P}}(\mathbf{X},U,(\mathbf{V}_{i})_{i})-\mathbf{X}$, follows the pdf $f$, independent of $\mathbf{X}$.
II-C Differential Privacy (DP)
Since clients do not directly transmit their data to the server, the original FL framework [1] provides a certain level of privacy. Nonetheless, a significant amount of information can still be inferred from the shared data, e.g., model parameters induced by gradient descent, by potential eavesdroppers within the FL network. Consequently, a privacy mechanism, such as DP, is essential to protect the shared information.
In this section, we delve into a notion of DP [10, 11]. In DP, clients place trust in the server (or data curator) responsible for collecting and holding their individual data in a database $X\in\mathcal{X}$, where $\mathcal{X}$ denotes the collection of databases. The server then introduces privacy-preserving noise to the original datasets or query results through a randomized mechanism $\mathcal{A}$, producing an output $Y=\mathcal{A}(X)\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of possible outputs, before sharing them with untrusted data analysts. While this model requires a higher level of trust compared to the local model, it enables the design of significantly more accurate algorithms. Two databases $X$ and $X^{\prime}$ are considered adjacent if they differ in only one entry. More generally, we can define a symmetric adjacent relation $\mathcal{R}\subseteq\mathcal{X}^{2}$ and say that $X$ and $X^{\prime}$ are adjacent databases if $(X,X^{\prime})\in\mathcal{R}$. Here, we review the definition of $(\epsilon,\delta)$-differentially private, initially introduced by Dwork et al. [10].
Definition 6 ($(\epsilon,\delta)$-differential privacy [10])
A randomized mechanism $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ with the associated conditional distribution $P_{Y|X}$ of $Y=\mathcal{F}(X)$ is $(\epsilon,\delta)$-differentially private ($(\epsilon,\delta)$-DP) if for all $\mathcal{S}\subseteq\mathcal{Y}$ and for all $(X,X^{\prime})\in\mathcal{R}$,
$\mathbb{P}(\mathcal{F}(X)\in\mathcal{S})\leq e^{\epsilon}\mathbb{P}(\mathcal{F}(X^{\prime})\in\mathcal{S})+\delta.$
When $\delta=0$, we say that $\mathcal{F}$ is $\epsilon$-differentially private ($\epsilon$-DP).
As the applications of differential privacy continue to expand, the issue of managing the privacy budget is gaining increasing attention, particularly focusing on privacy composition and privacy amplification. Privacy composition [11, 53, 54] asserts that the privacy budgets of composition blocks accumulate, while privacy amplification [38, 39] provides tools to analyze and bound the privacy budget of a combination of a selected base privacy mechanism to be less than the privacy of its constituent parts.
For our purposes, it is instrumental to express differential privacy in terms of $e^{\epsilon}$-divergence [55]. Let $\mathfrak{M}(\mathcal{Y})$ denote the set of probability measures on the output space $Y$. The $e^{\epsilon}$-divergence between two probability measures $\mu,\mu^{\prime}\in\mathfrak{M}(\mathcal{Y})$ is defined as:
$D_{e^{\epsilon}}(\mu||\mu^{\prime}):=\sup_{\mathcal{S}:\mathrm{\;measurable\;}\mathcal{S}\subseteq\mathcal{Y}}(\mu(\mathcal{S})-e^{\epsilon}\mu^{\prime}(\mathcal{S})).$ (9)
Note that $\mathcal{F}$ is $(\epsilon,\delta)$-differential privacy if and only if $D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime}))\leq\delta$ for every $(X,X^{\prime})\in\mathcal{R}$.
To study the relavant properties of $\mathcal{F}$ from a privacy amplification point of view, we review the notions of privacy profile and group privacy profile [38] as follows. The privacy profile $\delta_{\mathcal{F}}$ of a mechanism $\mathcal{F}$ is a function associating to each privacy parameter $e^{\epsilon}$ a supremium on the $e^{\epsilon}$-divergence between the outputs of running the mechanism on two adjacent databases, i.e.,
$\delta_{\mathcal{F}}(\epsilon):=\sup_{(X,X^{\prime})\in\mathcal{R}}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})).$ (10)
The group privacy profile $\delta_{\mathcal{F},j}(\tilde{\epsilon})$ ($j\geq 1$) is defined as:
$\delta_{\mathcal{F},j}(\tilde{\epsilon}):=\sup_{d(X,X^{\prime})\leq k}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})),$ (11)
where $d(X,X^{\prime}):=\min\{k:\exists X_{1},\ldots,X_{k-1}\mathrm{\;with\;}(X,X_{1}),\ldots,(X_{k-1},X^{\prime})\in\mathcal{R}\}$. Note that $\delta_{\mathcal{F}}=\delta_{\mathcal{F},1}$.
In [38], Balle et al. demonstrated a result on privacy amplification with subsampling with replacement as follows.
Theorem 7
[38, Theorem 10] Let $s$ be the subsampling with replacement mechanism and let $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ be a base privacy mechanism. Then, for $\tilde{\epsilon}\geq 0$ and set $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$, we have
$\delta_{\mathcal{F}(s(X))}(\epsilon)\leq\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\delta_{\mathcal{F}(X),j}(\tilde{\epsilon}),$ (12)
where $\delta_{\mathcal{F}(\cdot)}(\cdot)$ is the privacy profile of the privacy mechanism $\mathcal{F}(\cdot)$ and $\delta_{\mathcal{F}(\cdot),\cdot}(\cdot)$ is the group privacy profile of $\mathcal{F}(\cdot)$.
The above result guarantees that utilizing a differentially private mechanism on a random subsample of a dataset in a single communication round of FL framework offers stronger privacy assurances compared to applying it on the whole dataset.
Recall that for a function $q:\mathcal{X}\to\mathbb{R}^{n}$ with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, the Gaussian mechanism $\mathcal{F}(X)=q(X)+\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ satisfies $(\epsilon,\delta)$-DP if $\tilde{\sigma}^{2}\geq 2\Delta_{2}^{2}\log(1.25/\delta)/\epsilon^{2}$ and $\epsilon\in(0,1)$ [11, Appendix A]. In [56], Balle and Wang gave a new analysis of the Gaussian mechanism that is valid for all $\epsilon>0$.
Theorem 8
[56, Theorem 8] Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the Gaussian mechanism $\mathcal{F}(X):=q(X)+\mathbf{Z}$ with $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, is $(\epsilon,\delta)$-DP if and only if
$\Phi\left(\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\epsilon\tilde{\sigma}}{\Delta_{2}}\right)-e^{\epsilon}\Phi\left(-\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\epsilon\tilde{\sigma}}{\Delta_{2}}\right)\leq\delta,$ (13)
where $\Phi$ denotes the Gaussian cdf.
The above result can be interpreted as providing an expression for the privacy profile of the Gaussian mechanism in terms of the cdf of a standard Gaussian distribution.
We prove a simple consequence of Theorem 7 and Theorem 8 as follows.
Lemma 9
Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, and let $s$ be the subsampling with replacement mechanism. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the subsampled Gaussian mechanism $M(X)=q\circ s(X)+N(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ is $(\epsilon,\delta)$-DP if and only if, for any $\tilde{\epsilon}>0$, $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$ and
$\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\left(\Phi\left(\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)-e^{\tilde{\epsilon}/j}\Phi\left(-\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)\right)\leq\delta.$ (14)
Proof:
A standard group privacy analysis [57] gives the privacy profile
$\delta_{\mathcal{F}(X),j}(\tilde{\epsilon})\leq\frac{(e^{\tilde{\epsilon}}-1)\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j)}{e^{\tilde{\epsilon}/j}-1}.$
Thus, the RHS of (12) in Theorem 7 is upper-bounded by
$\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j)}{e^{\tilde{\epsilon}/j}-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\left(\Phi\left(\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)-e^{\tilde{\epsilon}/j}\Phi\left(-\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)\right),$
where (a) holds since the base mechanism $\mathcal{F}(\cdot)$ is Gaussian, which has the privacy profile given as the LHS of (13). This completes the proof. ∎
Another important privacy mechanism is the Laplace mechanism. The privacy profile of the Laplace mechanism is given as follows:
Theorem 10
[38, Theorem 3] Let $q:\mathcal{X}\rightarrow\mathbb{R}$ be a function with the global $\ell_{1}$-sensitivity $\Delta_{1}:=\sup_{(X,X^{\prime})\in\mathcal{R}}|q(X)-q(X^{\prime})|$ and $\mathcal{F}(X):=q(X)+Z$ with $Z\sim\mathrm{Lap}(0,b)$ be the Laplace mechanism. Then, for any $\epsilon>0$, the privacy profile of $\mathcal{F}$ is given by $\delta_{\mathcal{F}}(\epsilon)=\left[1-e^{\frac{\epsilon-\Delta_{1}/b}{2}}\right]_{+}$, where $[\cdot]_{+}:=\max\{\cdot,0\}$.
Note that $\delta_{\mathcal{F}}(\epsilon)=0$ for any $\epsilon\geq\Delta_{1}/b$, which implies the well-known fact that the Laplace mechanism with $b\geq\Delta_{1}/\epsilon$ is $\epsilon$-DP.
II-D Problem Setting
Our goal is to construct a mechanism that simultaneously addresses privacy requirements (protecting against honest but inquisitive clients within the FL network or inquisitive data analysts upon public release of the final trained model, see Section II-D1) and compression demands (for lossless uplink channels with limited bandwidth) in a single local model update at each client within the FL framework. Since the distribution of the model parameters and/or the induced gradient is often unknown to the clients, our focus lies in creating a universal mechanism applicable to any random source. Furthermore, we are exploring privacy mechanisms that introduce noise customized to various noise distributions. While Laplace mechanism provides pure DP protection, Gaussian mechanism only provides approximate DP protection with a small failure probability [10, 11]. However, it is well-known that Gaussian mechanism support tractability of the privacy budget in mean estimation [58, 54], an important subroutine in FL. Such schemes can be formulated as mappings from the local update $\mathbf{X}_{t}^{k}\in\mathbb{R}^{m}$ at client $k$ at the time instance $t$ to the estimated update $\hat{\mathbf{X}}_{t}^{k}\in\mathbb{R}^{m}$ at the server, aimed to achieve the following desired properties:
  1. 1.
    Privacy requirement : The perturbed query function generated by the privacy mechanism must adhere to $(\epsilon,\delta)$-DP (or $\epsilon$-DP). For instance, ensuring that the mapping of the average of local model updates across clients $\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbf{X}_{t}^{k}$ to the average of estimated updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$ at the server satisfies $(\epsilon,\delta)$-DP.
  2. 2.
    Compression/communication-efficiency : The estimation $\hat{\mathbf{X}}_{t+\tau}^{k}$ from client $k$ to the server should be represented by finite bits per sample.
  3. 3.
    Universal source : the scheme should operate reliably irrespective of the distribution of $\mathbf{X}_{t}^{k}$ and without prior knowledge of it.
  4. 4.
    Adaptable noise : The noise $Z$ in the privacy mechanism is customizable according to the required accuracy level and privacy protection.
II-D1 Threat model
We adopt trusted aggregator model, i.e., the server is trusted. Additionally, we assume that there are separate sources of shared randomness between each client and the server to perform quantization. Clients engaged in the FL framework are assumed to be honest yet inquisitive, meaning they comply with the protocol but may attempt to deduce sensitive client information from the average updates received by the server. Our goals are:
  • Protecting the privacy of each client’s local dataset from other clients, as the updated model between rounds may inadvertently disclose sensitive information.
  • Preventing privacy leaks from the final trained model upon completion of training, as it too may inadvertently reveal sensitive information. This ensures the relevance of our solution in scenarios where clients are trusted, and the final trained model may be publicly released to third parties.
III CEPAM for FL
In this section, we introduce CEPAM, which utilizes the inherent randomness in probabilistic quantized FL to provide privacy enhancement. (see Algorithm 1). Our approach modifies the schemes in [9] and [22] by replacing universal quantization [46, 47] with RSUQ as outlined in Section II-B. Through the utilization of RSUQ and the layered construction [35, 34], leveraging on the unique quantization error property of RSUQ (refer to Proposition 5), we construct a joint privacy and compression mechanism to address both FL challenges of communication overload and privacy requirement adaptable to any system adjustment simultaneously. This is achieved by applying LRSUQ on the model updates $\mathbf{X}_{t+\tau}^{k}$ at the end of each FL round for every client separately. These quantized updates are then transmitted as a set of messages $\{\mathbf{M}_{j}^{k}\}_{j\in\mathcal{N}}$ to the server. The server adds the dithers induced by the shared randomness to the received messages, collects the decoded messages and obtains the estimations $\hat{\mathbf{X}}_{t}^{k}$. Subsequently, the server computes the average of estimated model updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$, which ensures the DP guarantee against the honest but inquisitive clients. The performance of CEPAM will be analyzed in Section IV.
Algorithm 1 CEPAM
1:Inputs: Number of total iterations $T$, number of local iterations $\tau$, number of clients $K$, local datasets $\{\mathcal{D}^{(k)}\}_{k\in\mathcal{K}}$, loss function $\ell(\cdot,\cdot)$
2:Output: Global optimized model $\mathbf{W}_{T}$
3:Initialization: Client $k$ and the server agree on privacy budget $\epsilon>0$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$-DP (or privacy budget $\epsilon>0$ for $\epsilon$-DP), shared seed $s_{k}$, lattice dimension $n$, generator matrix $\mathbf{G}$, scaling coefficient $\gamma>0$, privacy-preserving noise $\mathbf{Z}\sim f$ with noise variance $\mathrm{Var}(f)>0$, latent variable $U\sim g(u):=\mu(L_{u}^{+}(f))$, initial model parameter vector $\mathbf{W}_{0}$
4:Protocol at client $k$:
5:for $t+1\notin\mathcal{T}_{T}$ do
6:     Receive $\mathbf{W}_{t}$ from the server or use $\mathbf{W}_{0}$ if $t=0$
7:     Set $\mathbf{W}_{t}^{k}\leftarrow\mathbf{W}_{t}$
8:     for $t^{\prime}=1$ to $\tau$ do
9:         Compute $\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t}^{k}}(\mathbf{W}_{t}^{k})$
10:     end for
11:     Compute $\mathbf{X}_{t+\tau}^{k}\leftarrow\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}$
12:     Run subroutine $\mathrm{ENCODE}(\mathbf{X}_{t+\tau}^{k},\gamma,s_{k},(f,g),N)$triangleright See Algorithm 2
13:     Send $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ to server, using $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits
14:end for
15:Protocol at the server:
16:for $t+\tau\in\mathcal{T}_{T}$ do
17:     for $k\in\mathcal{K}$ do
18:         Run subroutine $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},\gamma,s_{k},g,N)$triangleright See Algorithm 3
19:     end for
20:     Compute $\hat{\mathbf{W}}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}$
21:     Set $\mathbf{W}_{t+\tau}\leftarrow\hat{\mathbf{W}}_{t+\tau}$ and broadcast $\mathbf{W}_{t+\tau}$ to all clients, or output $\mathbf{W}_{T}$ if $t+\tau=T$
22:end for
CEPAM consists of three stages: initialization, encoding at client, and decoding at server. The initialization stage involving both the clients and the server takes place before the start of the FL procedure, while the encoding and decoding stages are independently carried out by the clients and the server. Each client is presumed to execute the same encoding function, ensuring consistency across all clients. Thus, we focus on the $k$-client, and the details of each stage are described as follows.
1) Initialization: At the outset, client $k$ and the server agree on the privacy budget $\epsilon$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$ DP (or $\epsilon$ for $\epsilon$-DP) in accordance with the system requirements, along with the parameters for LRSUQ as specified in Section II-B. The latter involves sharing a small random seed $s_{k}$ between client $k$ and the server to serve as a source of common randomness while fixing the lattice dimension $n$, a lattice generator matrix $\mathbf{G}$, and a scaling coefficient $\gamma>0$. Both client $k$ and the server use the same random seed to initialize their respective PRNGs $\mathfrak{P}$ and $\tilde{\mathfrak{P}}$, ensuring that the outputs of the two PRNGs remain identical.
2) Client: At the end of each FL round of local training, when the model update $\mathbf{X}_{t+\tau}^{k}\in\mathbb{R}^{m}$ is ready for transmission through the binary uplink channel to the server, client $k$ executes the encoding process. The update $\mathbf{X}_{t+\tau}^{k}$ is encoded into finite bit representations by a combination of LRSUQ and the entropy coding, which also jointly ensures the privacy guarantee with the decoding step at the server. The encoding algorithm is summarized in Algorithm 2.
Algorithm 2 $\mathrm{ENCODE}(\mathbf{X}_{t+\tau}^{k},\gamma,s_{k},(f,g),N)$
1:Inputs: model update vector ${\mathbf{X}}_{t+\tau}^{k}$, scaling coefficient $\gamma$, random seed $s_{k}$, pair of pdfs $(f,g)$, number of sub-vectors $N$
2:Output: Set of messages $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$
3:Compute $\tilde{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{X}_{t+\tau}^{k}/(\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|)$triangleright Perform normalizing
4:Partition $\tilde{\mathbf{X}}_{t+\tau}^{k}$ to $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$
5:for $j=1$ to $N$ do
6:     Sample $U_{t+\tau,j}^{k}\sim g$ by $\mathfrak{B}^{k}$triangleright Initiated by seed $s_{k}$
7:     for $i=1,2,\ldots$ do $\triangleright$ Perform RSUQ
8:          Sample $\mathbf{V}_{t+\tau,j,i}^{k}\sim\mathrm{Unif}(\mathcal{P})$ by $\mathfrak{B}^{k}$
9:         Find unique $\mathbf{M}_{t+\tau,j}^{k}\leftarrow Q_{\mathcal{P}}(\tilde{\mathbf{X}}_{t+\tau,j}^{k}/\beta(U_{t+\tau,j}^{k})-\mathbf{V}_{t+\tau,j,i}^{k})\in\mathbf{G}\mathbb{Z}^{n}$triangleright Perform encoding
10:          Check if $\beta(U_{t+\tau,j}^{k})\cdot\mathbf{M}_{t+\tau,j}^{k}-\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in L_{U_{t+\tau,j}^{k}}^{+}(f)$
11:         if Yes then $\triangleright$ Perform rejection sampling
12:              $H_{t+\tau,j}^{k}\leftarrow i$; Return $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})$
13:         else
14:              Reject $i$ and repeat Step 8-10 with $i+1$
15:         end if
16:     end for
17:end for
18:return $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$
Quantization: The $k$-th client normalizes the update $\mathbf{X}_{t+\tau}^{k}$ by a real number $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|>0$ to $\tilde{\mathbf{X}}_{t+\tau}^{k}$, and then divides $\tilde{\mathbf{X}}_{t+\tau}^{k}$ into $N:=\left\lceil\frac{m}{n}\right\rceil$ distinct $n$-dimensional sub-vectors $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathbb{R}^{n}$, $j\in\{1,\ldots,N\}=:\mathcal{N}$. The overhead associated with compressing the scalar $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|$ is assumed to be negligible compared to compressing the vector $\tilde{\mathbf{X}}_{t+\tau}^{k}$.
Instead of using the universal quantization, we replace it by LRSUQ (see Definition 4). Let $\mathbf{Z}\sim f$ denote the the intended privacy-preserving noise random vector with mean $\boldsymbol{\mu}:=\mathbb{E}[\mathbf{Z}]=\mathbf{0}$ and noise variance $\mathrm{Var}(f):=\mathrm{Var}(\mathbf{Z})$, and let $g(u):=\mu(L_{u}^{+}(f))$ be the pdf of some latent variable. To carry out LRSUQ at client $k$, the encoder at the client observes for $j\in\mathcal{N}$, $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathbb{R}^{n}$, generates $\mathbf{V}_{t+\tau,j,1}^{k},\mathbf{V}_{t+\tau,j,2}^{k},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ and $U_{t+\tau,j}^{k}\sim g$ by the PRNG $\mathfrak{P}^{k}$ until iteration $H_{t+\tau,j}^{k}$ satisfying $\beta(U_{t+\tau,j}^{k})\cdot\big{(}Q_{\mathcal{P}}(\tilde{\mathbf{X}}_{t+\tau,j}^{k}/\beta(U_{t+\tau,j}^{k})-\mathbf{V}_{H_{t+\tau,j}^{k}})+\mathbf{V}_{H_{t+\tau,j}^{k}}\big{)}-\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in L_{U_{t+\tau,j}^{k}}^{+}(f)$, computes $\mathbf{M}_{t+\tau,j}^{k}:=Q_{\mathcal{P}}(\tilde{\mathbf{X}}_{t+\tau,j}^{k}/\beta(U_{t+\tau,j}^{k})-\mathbf{V}_{H_{t+\tau,j}^{k}})\in\mathbf{G}\mathbb{Z}^{n}$, and encodes and transmits $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})$. [^3] By setting $U_{t+\tau,j}^{k}=\emptyset$, we could retrieve the simpler RSUQ as defined in Definition 2.
3) Server: The server observes $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ from client $k$. Using the shared randomness between client $k$ and the server, i.e., the shared random seed $s_{k}$, the same realizations of $\mathbf{V}_{H_{t+\tau,j}^{k}}$’s and $U_{t+\tau,j}^{k}$’s generated by client $k$ can also be obtained by the server. More specifically, the decoder generates $U_{t+\tau,j}^{k}\sim g$ and $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ by the PRNG $\tilde{\mathfrak{P}}^{k}$, and outputs $\mathbf{Y}_{t+\tau,j}^{k}=\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$. Subsequently, the decoder collects the sub-vectors $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into a $m\times 1$ and scales it by $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|$ to obtain an estimation $\hat{\mathbf{X}}_{t+\tau}^{k}$ of the update $\mathbf{X}_{t+\tau}^{k}$. The decoding algorithm is summarized in Algorithm 3.
Algorithm 3 $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},\gamma,s_{k},g,N)$
1:Inputs: Set of messages $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$, scaling coefficient $\gamma$, random seed $s_{k}$, pdf $g$, number of sub-vectors $N$
2:Output: Estimated model update $\hat{\mathbf{X}}_{t+\tau}^{k}$
3:Use $\tilde{\mathfrak{P}}^{k}$ to sample $U_{t+\tau,j}^{k}\sim g$triangleright Initiated by seed $s_{k}$
4:for $j\in\mathcal{N}$ do
5:     Receive $H_{t+\tau,j}^{k}$ and $\mathbf{M}_{t+\tau,j}^{k}$
6:     Use $\tilde{\mathfrak{P}}^{k}$ to sample $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$triangleright Initiated by seed $s_{k}$
7:      Compute $\mathbf{Y}_{t+\tau,j}^{k}\leftarrow\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$triangleright Perform decoding, inducing privacy protection
8:end for
9:Collect $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into $\mathbf{Y}_{t+\tau}^{k}$
10:Estimate $\hat{\mathbf{X}}_{t+\tau}^{k}\leftarrow\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|\cdot\mathbf{Y}_{t+\tau}^{k}$
11:return $\hat{\mathbf{X}}_{t+\tau}^{k}$
Privacy enhancement: The privacy is ensured by LRSUQ executed cooperatively between the clients and the server in compliance with the threat model and the defined privacy guarantee. Through the decoding process, it is ensured that $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ comprises noisy estimates of $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$, i.e., the estimated model updates at the server are noisy estimates of the model updates at the clients, thus establishing a privacy mechanism. Refer to Step 7 in Algorithm 3. Specifically, in Section IV-A1, we prove that when using the Gaussian mechanism, the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP requirement by specifying appropriate $f$ and $g$, while in Section IV-A2, when using the Laplace mechanism, the estimated model updates satisfies $\epsilon$-DP requirement by specifying appropriate $f$ and $g$.
IV Performance Analysis
In this section, we study the performance of CEPAM. This includes characterizing its privacy guarantees and compression capabilities, followed by an exploration of its distortion bounds and convergence analysis. We begin by stating a useful lemma for the subsequent analysis. The proof is given in Appendix -A.
Lemma 11
The LRSUQ quantization errors $\{\tilde{\mathbf{Z}}_{t+\tau,j}^{k}:=\mathbf{Y}_{t+\tau,j}^{k}-\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ are iid (over $k$ and $j$), follows the pdf $f$, and independent of $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$.
IV-A Privacy
By using LRSUQ, we can construct privacy mechanisms that satisfy various privacy requirements by customizing the pair $(f,g)$. In the following subsections, we present examples of $(f,g)$ pairs to illustrate the construction of Gaussian and Laplace mechanisms, along with demonstrating the associated privacy guarantees. For simplicity, we assume that $p_{k}=1/K$.
IV-A1 Gaussian mechanism
If $g$ follows a chi-squared distribution with $n+2$ degrees of freedom, i.e., $g\sim\chi_{n+2}^{2}$, then $f$ follows a Gaussian distribution. This guarantees that the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP. The proof is given in Appendix -B.
Theorem 12
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for every $k,j$, the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian. The average $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the average of the model updates such that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}=\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbf{X}_{t+\tau}^{k}+\mathcal{N}\bigg{(}\mathbf{0},\frac{\sigma^{2}\|\mathbf{X}_{t+\tau}^{k}\|^{2}}{K\gamma^{2}}\mathbf{I}_{m}\bigg{)},$ (15)
where $\mathbf{I}_{m}$ is the $m\times m$ identity matrix. Hence, for every $\tilde{\epsilon}>0$, $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ satisfies $(\epsilon,\delta)$-DP in one round against clients for $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p=1-\left(1-\frac{1}{|\mathcal{D}^{(k)}|}\right)^{\tau}$ and
$\delta=\sum_{j=1}^{\tau}\binom{\tau}{j}\bigg{(}\frac{1}{|\mathcal{D}^{(k)}|}\bigg{)}^{j}\bigg{(}1-\frac{1}{|\mathcal{D}^{(k)}|}\bigg{)}^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\left(\Phi\bigg{(}\frac{\tau\gamma}{\sqrt{K}\sigma}-\frac{\sqrt{K}\tilde{\epsilon}\sigma}{2j\tau\gamma}\bigg{)}-e^{\tilde{\epsilon}/j}\Phi\bigg{(}-\frac{\tau\gamma}{\sqrt{K}\sigma}-\frac{\sqrt{K}\tilde{\epsilon}\sigma}{2j\tau\gamma}\bigg{)}\right).$ (16)
IV-A2 Laplace mechanism
If $g$ follows a Gamma distribution $\mathrm{Gamma}(2,1)$, then $f$ follows a Laplace distribution.444 Our CEPAM can be extended to include multivariate Laplace distributions [59] or $t$-distributions [60], which are left for future work. This guarantees that estimated model updates satisfy $\epsilon$-DP. The proof is given in Appendix -C.
Theorem 13
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\mathrm{Gamma}(2,1)$ and $\tilde{Z}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}((-bu,bu))$ for every $k$ and $j$, the resulting mechanism is Laplace. The estimator $\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the model update $\mathbf{X}_{t+\tau}^{k}$ such that
$\hat{\mathbf{X}}_{t+\tau}^{k}=\mathbf{X}_{t+\tau}^{k}+\mathrm{Lap}\left(\mathbf{0},(b\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|)\mathbf{I}_{m}\right),$ (17)
where $\mathbf{I}_{m}$ is the $m\times m$ identity matrix. Hence, for every $\hat{\mathbf{X}}_{t+\tau}^{k}$ satisfies $(\epsilon,0)$-DP in one round against clients for $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p=1-\left(1-\frac{1}{|\mathcal{D}^{(k)}|}\right)^{\tau}$, and $\delta=0$ provided that $\tilde{\epsilon}\geq 2\tau\gamma/b$.
IV-B Compression
Every client in CEPAM is required to transmit the set of message pairs $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})_{j\in\mathcal{N}}$ per communication round. If we know that $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathcal{X}$, then we can compress $H_{t+\tau,j}^{k}$ using the optimal prefix-free code [44, 61] for $\mathrm{Geom}(p(u))|\{U_{t+\tau,j}^{k}=u\}$, where $p(u):=\mu(L_{u}^{+}(f))/\mu(\beta(u)\mathcal{P})$, and compress $\mathbf{M}_{t+\tau,j}^{k}|\{U_{t+\tau,j}^{k}=u\}\in\mathcal{M}:=(\mathcal{X}+L_{u}^{+}(f)-\beta(u)\mathcal{P})\cap\mathbf{G}\mathbb{Z}^{n}$ using $H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)$ bits. Therefore, the total communication cost per communication round per client is at most $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits.
IV-C Distortion bounds
Representing the clipped model update $\tilde{\mathbf{X}}_{t+\tau}^{k}$ of client $k$ using a finite number of bits inherently introduces distortion, i.e., the recovered vector $\hat{\mathbf{X}}_{t+\tau}^{k}=\tilde{\mathbf{X}}_{t+\tau}^{k}+\mathbf{Z}_{t+\tau}^{k}$. We can then characterize the squared norm of the FL quantization error. The proof is given in Appendix -D.
Proposition 14
The FL quantization error $\mathbf{Z}_{t+\tau}^{k}$ has zero mean and satisfies
$\mathbb{E}[\|\mathbf{Z}_{t+\tau}^{k}\|^{2}\big{|}\mathbf{x}_{t+\tau}^{k}]=\gamma^{-2}\|\mathbf{x}_{t+\tau}^{k}\|^{2}N\mathrm{Var}(f).$ (18)
At the end of each FL round, CEPAM incorporates a DP mechanism based on LRSUQ, inherently introducing distortion as mentioned above. This distortion emerges as the update $\mathbf{X}_{t+\tau}^{k}$ is mapped into its distorted counterpart $\hat{\mathbf{X}}_{t+\tau}^{k}$. Consequently, the global model (4) becomes
$\hat{\mathbf{W}}_{t+\tau}:=\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}=\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{W}}_{t+\tau}^{k},$ (19)
where $\hat{\mathbf{W}}_{t+\tau}^{k}=\mathbf{W}_{t}+\hat{\mathbf{X}}_{t+\tau}^{k}$. Under common assumptions used in FL analysis, we can bound the error between $\mathbf{W}_{t+\tau}$ and $\hat{\mathbf{W}}_{t+\tau}$. To that end, we adopt the following assumptions:
AS1: The $\xi_{k,j}$’s, where $j=1,\ldots,n_{k}$, are i.i.d. samples in $\mathcal{D}^{(k)}$. However, different datasets $\mathcal{D}^{(k)}$ can be statistically heterogeneous and follow different distributions.
AS2: Let the sample index $j_{t}^{k}$ be sampled from client $k$’s local data $\mathcal{D}^{(k)}$ uniformly at random. The expected squared norm of stochastic gradients is bounded by some $\theta_{k}^{2}>0$, i.e., $\mathbb{E}[\|\nabla F_{k}^{j_{t}^{k}}(\mathbf{W})\|]\leq\theta_{k}^{2}$, for all $\mathbf{W}\in\mathbb{R}^{m}$.
We have the following bound on the error between $\mathbf{W}_{t+\tau}$ and $\hat{\mathbf{W}}_{t+\tau}$. The proof is given in Appendix -E.
Proposition 15
When AS2 holds, the mean-squared error $\mathbb{E}[\|\hat{\mathbf{W}}_{t+\tau}-\mathbf{W}_{t+\tau}\|^{2}]$ is upper-bounded by
$N\gamma^{-2}\tau\mathrm{Var}(f)\sum_{t^{\prime}=t}^{t+\tau-1}\eta_{t^{\prime}}^{2}\sum_{k\in\mathcal{K}}p_{k}^{2}\theta_{k}^{2}.$ (20)
IV-D FL Convergence analysis
We next study the FL convergence of CEPAM, considering the following common assumptions used in the FL [40, 41, 9]:
AS3: $F_{1},\ldots,F_{K}$ are all $L$-smooth, i.e., for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{m}$, $F_{k}(\mathbf{y})-F_{k}(\mathbf{x})\leq(\mathbf{y}-\mathbf{x})^{T}\nabla F_{k}(\mathbf{x})+\frac{L}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$.
AS4: $F_{1},\ldots,F_{K}$ are all $C$-strongly convex, [^5] The extension to non-convex objective functions is left for future work. i.e., for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{m}$, $F_{k}(\mathbf{y})-F_{k}(\mathbf{x})\geq(\mathbf{y}-\mathbf{x})^{T}\nabla F_{k}(\mathbf{x})+\frac{C}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$.
We define the following to quantify the degree of non-iid:
$\psi=F(\mathbf{w}^{*})-\sum_{k\in\mathcal{K}}\min_{\mathbf{w}\in\mathbb{R}^{m}}F_{k}(\mathbf{w}),$ (21)
where $\mathbf{w}^{*}$ denotes the minimum of $F(\cdot)$. If the data is iid, meaning that the training data originates from the same distribution, then $\psi$ goes to zero as the training size increases. In the case of non-i.i.d. data, $\psi$ is positive, and its magnitude reflects the heterogeneity of the data distribution.
The convergence of CEPAM for FL learning is given as follows. The proof is given in Appendix -F.
Theorem 16
Assume that AS1-4 hold and $\theta_{k}$ $L$, $C$ be defined therein. Set $\alpha=\tau\max\{\frac{4L}{C},1\}$ and the learning rate $\eta_{t}=\frac{\zeta}{t+\alpha}$ for nonnegative integers $t\geq 0$. Then, CEPAM satisfies
$\mathbb{E}[F(\mathbf{W}_{T})]-F(\mathbf{w}^{*})\leq\frac{L}{2(T+\alpha)}\max\left\{\frac{B\tau^{2}}{C^{2}(\tau-1)},\alpha\|\mathbf{W}_{0}-\mathbf{w}^{*}\|\right\},$ (22)
where
$B:=6L\psi+\left(4N\gamma^{-2}\tau^{2}\mathrm{Var}(f)+1\right)\sum_{k\in\mathcal{K}}p_{k}^{2}\theta_{k}^{2}+8(\tau-1)^{2}\sum_{k\in\mathcal{K}}p_{k}\theta_{k}^{2}.$
V Numerical Evaluations
In this section, we evaluate the performance of CEPAM. [^6] The source code used in our numerical evaluations is available at https://github.com/yokiwuuu/CEPAM.git. We begin by detailing our experimental setup, which includes the types of datasets, model architectures, and training configurations in Section V-A. Afterwards, we present comprehensive experimental results to demonstrate the effectiveness of CEPAM by comparing it to several baselines realizing the privacy protection and quantization in the FL framework in Section V-B.
We also compare the accuracy-privacy trade-off between CEPAM and the Gaussian-mechanism-then-quantize approach in Section V-C. All experiments were executed on a server equipped with dual Intel(R) Xeon(R) Platinum 8480C CPUs (112 cores in total), 512GB RAM, and an NVIDIA RTX A5000 GPU (24GB VRAM) running Ubuntu 22.04.4 LTS. The implementation was based on PyTorch and CUDA 12.5 with NVIDIA driver version 555.42.06.
V-A Experimental Setup
V-A1 Datasets
We evaluate CEPAM on the standard image classification benchmark MNIST, which consists of $28\times 28$ grayscale handwritten digits images divided into 60,000 training examples and 10,000 test examples. The training examples and the test examples are equally distributed among $K=30$ clients. For simplicity, we set $p_{k}=1/K$ for $k\in\mathcal{K}$.
V-A2 Learning Architecture
We evaluate CEPAM using two different neural network structures: a multi-layer perceptron (MLP) with two hidden layers and intermediate ReLU activations; and a convolutional neural network (CNN) composed of two convolutional layers followed by two fully-connected ones, with intermediate ReLU activations and max-pooling layers. All two models use a softmax output layer. There are 6422 learnable parameters for CNN and 25818 learnable parameters for MLP.
V-A3 Baselines
We compare CEPAM to other baselines:
  • FL: vanilla FL without any privacy or compression.
  • FL+SDQ : FL with scalar SDQ-based compression, no privacy.
  • FL+{Gaussian, Laplace}: FL with one-dimensional Gaussian or Laplace mechanism, no compression
  • FL+{Gaussian, Laplace}+SDQ: FL with approach that applies one-dimensional Gaussian or Laplace mechanism followed by scalar SDQ.
  • CEPAM-{Gaussian, Laplace}: CEPAM achieves privacy and quantization jointly through LRSUQ to construct Gaussian or Laplace mechanisms, namely CEPAM-Gaussian or CEPAM-Laplace. In particular, for CEPAM-Gaussian, we evaluate three different cases of LRSUQ that simulate Gaussian noise for dimension $n=1,2,3$, where we use integer lattice $\mathbb{Z}^{n}$, [^7] To simplify implementation, we choose to use the integer lattice due to its well-established decoding algorithm [62]. The time complexity of the rejection sampling step can be improved by using a lattice with higher packing density. and the corresponding basic cells $(-0.5,0.5]$, $(-0.5,0.5]^{2}$, and $(-0.5,0.5]^{3}$. For CEPAM-Laplace using LRSUQ to simulate Laplace noise, we use integer lattice $\mathbb{Z}$ with the basic cell $(-0.5,0.5]$.
V-A4 Training Configurations
We select the momentum SGD as the optimizer, where the momentum is set to $0.9$. The local iterations $\tau$ per communication round is set to 15. While we set the initial learning rate to be $0.01$, we also implement an adaptive learning rate scheme that reduces the learning rate by a factor of $0.5$ when the validation accuracy plateaus for $10$ consecutive rounds.
V-B FL Convergence
We report the FL convergence of CEPAM-Gaussian and CEPAM-Laplace in terms of accuracy using MLP and CNN architectures and MNIST dataset.
V-B1 Gaussian Mechanism
We set $\sigma=0.1$ and $\tilde{\epsilon}=5.9$ for the base Gaussian mechanisms in CEPAM-Gaussian for $n=1,2,3$ with variances $0.1^{2}$, $2\times 0.1^{2}$, $3\times 0.1^{2}$, respectively, and all the clients’ composite Gaussian mechanism achieve $(\epsilon=1.45,\delta=9.69\times 10^{-3})$-DP. We also set $\sigma=0.1$ for both FL+Gaussian and FL+Gaussian+SDQ.
Figure 1 shows the validation accuracy of CEPAM-Gaussian over communication rounds using MLP and CNN model, respectively. Note that for MLP, the cases of CEPAM-Gaussian ($n=1,2,3$) consistently outperform and achieve higher accuracy than all the baselines. Similar trend can be observed for CNN.
(a) Convergence profiles of different FL schemes using MLP.
(b) Convergence profiles of different FL schemes using CNN.
Figure 1: Validation accuracy curves of CEPAM-Gaussian against Baselines using different neural network models.
In Table I, we report the test accuracy of CEPAM-Gaussian for different architectures (MLP and CNN) using MNIST. CEPAM-Gaussian demonstrates better performance, achieving an improvement of 1.0-2.0% and 0.5-1.0% in accuracy compared to other baselines for MLP and CNN, respectively, suggesting that CEPAM can be beneficial for various learning models. Furthermore, all cases of CEPAM-Gaussian have approximately the same accuracy. This is because the added noise per dimension is consistent, i.e., $\mathrm{Var}(f)/n=\sigma^{2}$. However, higher-dimensional LRSUQ has a better compression ratio than lower-dimensional counterparts due to its nature as a vector quantizer. In contrast, for baselines such as the Gaussian-mechanism-then-quantize approach (Gaussian+SDQ), the total noise comprises privacy noise and quantization error, leading to more distortion and consequently lower accuracy compared to LRSUQ. It is worth noting that adding a minor level of distortion as a regularizer during training deep models can potentially enhance the performance of the converged model [63]. This finding is corroborated in our experimental results, where CEPAM-Gaussian achieves slightly better accuracy than vanilla FL for both architectures.
Table I: Test Accuracy (Gaussian) for MNIST
Baselines MLP (%) CNN (%)
FL 82.50 91.34
FL+SDQ 82.42 91.46
FL+Gaussian 83.14 91.39
FL+Gaussian+SDQ 83.15 90.74
CEPAM-Gaussian ($n=1$) 84.42 91.83
CEPAM-Gaussian ($n=2$) 84.53 91.61
CEPAM-Gaussian ($n=3$) 83.85 91.94
V-B2 Laplace Mechanism
We set $b=0.1$ and $\tilde{\epsilon}=5000$ for the base Laplace mechanism in CEPAM-Laplace with a variance of $2\times 0.1^{2}$, and the composite Laplace mechanism achieves $(\epsilon=4995)$-DP. We also set $b=0.1$ for both FL+Laplace and FL+Laplace+SDQ. Note that a higher privacy budget is required to ensure $\epsilon$-DP for CEPAM-Laplace under privacy amplification with subsampling analysis. However, we might achieve a lower privacy budget by relaxing $\delta$. The trade-off is left for the future study.
Figure 2 shows the validation accuracy of CEPAM-Laplace over communication rounds using the MLP and CNN models, respectively. Similar to CEPAM-Gaussian, we observe that CEPAM-Laplace consistently outperforms other baselines.
(a) Convergence profiles of different FL schemes using MLP.
(b) Convergence profiles of different FL schemes using CNN.
Figure 2: Validation accuracy curves of CEPAM-Laplace against Baselines using different neural network models.
In Table II, we also report the performance of test accuracy of CEPAM-Laplace. We observe that CEPAM-Laplace achieves an improvement of about 2.0-2.7% and 1.5-2.0% in accuracy compared to other baselines for MLP and CNN, respectively, which suggests that CEPAM-Laplace can be beneficial for many learning models.
Table II: Test Accuracy (Laplace) for MNIST
Baselines MLP (%) CNN (%)
FL 83.02 91.08
FL+SDQ 83.28 91.00
FL+Laplace 83.78 90.59
FL+Laplace+SDQ 83.78 90.86
CEPAM-Laplace 85.76 92.66
V-C Privacy-Accuracy Trade-off
Figure 3 illustrates the trade-off between learning performance, measured in terms of test accuracy, and the privacy budget between CEPAM and the simple Gaussian-mechanism-then-quantize approach (Gaussian+SDQ) using a CNN architecture. Overall, CEPAM-Gaussian outperforms Gaussian+SDQ. The figure demonstrates that as more privacy budget is allocated, higher test accuracy can be achieved with CEPAM-Gaussian. However, there is a point of diminishing returns; once the privacy budget reaches a threshold (i.e., $\epsilon=3.0$), the increase in test accuracy by further increasing the privacy budget becomes limited.
Figure 3: Learning Accuracy and Privacy Trade-off using CNN
VI Conclusion and Discussion
In this paper, we introduced CEPAM, which can be used to achieved communication efficiency and privacy protection simultaneously in FL framework. We also proved the FL convergence bound for CEPAM and provided a privacy amplication analysis for CEPAM-Gaussiang and CEPAM-Laplace. Furthermore, the experimental results confirmed our theoretical findings through a comparison of CEPAM with alternative methods using the MNIST dataset with MLP and CNN architectures. The experimental outcomes demonstrated that CEPAM-Gaussian and CEPAM-Laplace lead to a moderate enhancement in accuracy by 0.5-2.0% and 1.5-2.7%, respectively, when contrasted with other baseline approaches.
For potential future directions, it may be of interest to investigate the performance of CEPAM for the non-convex smooth objective function. Moreover, CEPAM could be expanded to incorporate other nonuniform continuous noise distributions, such as $t$-distributions or multivariate Laplace distributions, in order to provide alternative privacy mechanisms.